# Degree-Weighted DeGroot Learning

### v2

@@ B.4.2 Proof of the Second Part of Theorem 1
$|\lambda_{2}(\mathbf{T})-\lambda_{2}(\mathbf{T}^{*})|\leq\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert,$ (36)
[…]
Under Assumptions (1) and (2), the expected degrees satisfy:
$1<\liminf_{n}\frac{\min_{i}d_{i}(\mathbf{R})}{\max_{i}d_{i}(\mathbf{R})}\leq\limsup_{n}\frac{\max_{i}d_{i}(\mathbf{R})}{\min_{i}d_{i}(\mathbf{R})}<\infty,$
which implies the existence of a positive constant $C$ such that for all $n$,
$\min_{i}d_{i}(\mathbf{R})<\max_{i}d_{i}(\mathbf{R})<C\min_{i}d_{i}(\mathbf{R}).$ (37)
Furthermore, we have
$\min_{i}(\mathbf{D}_{\mathbf{R}})_{ii}=\min_{i}\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}=\min_{i}\sum_{j}R_{ij}\frac{\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}\geq\min_{i}d_{i}(\mathbf{R}).$ (38)
First, note that the spectral norm in (36) can be decomposed as:
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert$ (39)
@@
$\|\mathbf{A}\|\leq\|\mathbf{A}-\mathbf{R}\|+\|\mathbf{R}\|\,.$ (40)
@@
$\|\mathbf{R}\|\leq\|\mathbf{R}\|_{F}=\sqrt{\sum_{i,j}R^{2}_{ij}}\leq Cn\max_{i,j}P_{ij}\leq C\min_{j}d_{j}(\mathbf{R})\,,$ (41)
@@
Applying Proposition 4 along with Assumption 1, we obtain
$\|\mathbf{A}\|\leq C\min_{j}d_{j}(\mathbf{R})\,,$ (42)
with probability at least $1-\frac{2}{n^{2}}$. Similarly, by Proposition 5,
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\right\rVert\leq C\left(\min_{j}d_{j}(\mathbf{R})\right)^{-1/2}\,,$ (43)
with probability at least $1-\frac{14}{n^{2}}$, for sufficiently large $n$.
@@
$\left\{\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\right\}\bigcap\left\{\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\right\},$ (44)
which holds with probability at least $1-\frac{16}{n^{2}}$, we derive
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert\leq C\frac{\sqrt{n\log n}}{\min_{i}d_{i}(\mathbf{R})}=C\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,.$ (45)
@@
B.5 Supplementary Proofs of Theorem 1
@@
To bound $\|\mathbf{A}-\mathbf{R}\|$, we rewrite it as:
$\mathbf{A}-\mathbf{R}=\mathbf{Y}+\sum_{1\leq i<j\leq n}\mathbf{X}^{i,j},$
@@
$\mathbf{X}^{i,j}=(\mathbf{A}-\mathbf{R})(\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}),$
$\mathbf{Y}=\operatorname{diag}(A_{11}-R_{11},\ldots,A_{nn}-R_{nn}),$
and $\mathbf{e}_{i}$ is the standard basis vector in $\mathbb{R}^{n}$.
Observe that $\mathbf{X}^{i,j}$ are independent, have mean zero, and are also independent of $\mathbf{Y}$, which itself has mean zero. Then, applying the matrix Bernstein inequality (Theorem 5.4.1 in Vershynin (2018)), we obtain:
$\operatorname{\mathbb{P}}\left\{\left\lVert\mathbf{Y}+\sum_{1\leq i<j\leq n}\mathbf{X}^{i,j}\right\rVert\geq t\right\}\leq 2n\exp\left(-\frac{t^{2}/2}{\sigma^{2}+Kt/3}\right),$ (46)
where
$\|\mathbf{Y}\|,\left\lVert\mathbf{X}^{i,j}\right\rVert\leq K=1,$
and
$\sigma^{2}=\left\lVert{\rm I}\kern-1.79993pt{\rm E}\mathbf{Y}^{2}+\sum_{1\leq i<j\leq n}{\rm I}\kern-1.79993pt{\rm E}(\mathbf{X}^{i,j})^{2}\right\rVert.$
Since
$(\mathbf{X}^{i,j})^{2}=(A_{ij}-R_{ij})^{2}(\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}),$
we obtain
${\rm I}\kern-1.79993pt{\rm E}(\mathbf{X}^{i,j})^{2}=(R_{ij}-R_{ij}^{2})(\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}).$
Similarly,
${\rm I}\kern-1.79993pt{\rm E}\mathbf{Y}^{2}=\sum_{i=1}^{n}(R_{ii}-R_{ii}^{2})\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
Since each $A_{ij}$ follows a Bernoulli distribution, we have $R_{ij}\in[0,1]$, and thus, $R_{ij}-R_{ij}^{2}\leq 1/4$. It follows that
$\sigma^{2}\leq n/4.$
Setting $t=3\sqrt{n\log n}$ in (46), we obtain:
$\operatorname{\mathbb{P}}\left\{\left\lVert\sum_{1\leq i<j\leq n}\mathbf{Y}+\mathbf{X}^{i,j}\right\rVert\geq 3\sqrt{n\log n}\right\}$ $\leq 2n\exp\left(-\frac{\frac{9n\log n}{2}}{\frac{n}{4}+\frac{3\sqrt{n\log n}}{3}}\right)$
$\leq 2n\exp(-3\log n)$
$=\frac{2}{n^{2}}.$
$\Box$
@@
The proof of Proposition 5 follows from Lemmas 4 to 6. Lemmas 4 and 5 have already been stated; we now proceed to prove them.
Proof of Lemma 4:
We first prove (27). The inequality
$\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))$
follows from the monotonicity property of $\phi$ (Property 1) for $\alpha\geq 0$. For the reverse inequality, we consider
$\log\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}$ $=\log\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))-\log\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))$
$=(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial}{\partial d}\log\phi(\alpha,d_{0}),$
for some $d_{0}\in[\min_{i}d_{i}(\mathbf{R}),K\min_{i}d_{i}(\mathbf{R})]$.
By Property 3 in (13),
$(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial}{\partial d}\log\phi(\alpha,d_{0})$ $=(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial\phi/\partial d(\alpha,d_{0})}{\phi(\alpha,d_{0})}$
$\leq(K-1)Kd_{0}\frac{\partial\phi/\partial d(\alpha,d_{0})}{\phi(\alpha,d_{0})}\leq C.$
Thus, we obtain
@@
$\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}=\exp\left(\log\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}\right)\leq e^{C}.$
The proof of (28) follows similarly using Property 3 in (14). Finally, (29) and (30) follow from (27) and (28), respectively, by reversing all inequalities, since $\partial\phi/\partial d\leq 0$ for $\alpha<0$.
$\blacksquare$
Proof of Lemma 5:
Recall that
$d_{j}({\mathbf{A}})=\sum_{i=1}^{n}A_{ij},\quad d_{j}(\mathbf{R})=\sum_{i=1}^{n}R_{ij}.$
Applying Hoeffding’s inequality, we obtain
$\operatorname{\mathbb{P}}\{|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\geq t\}$ =$=\operatorname{\mathbb{P}}\left\{\left|\sum_{i=1}^{n}(A_{ij}-R_{ij})\right|\geq t\right\}$
$\leq\operatorname{\mathbb{P}}\left\{\sum_{i=1}^{n}(A_{ij}-R_{ij})\geq t\right\}+\operatorname{\mathbb{P}}\left\{\sum_{i=1}^{n}(R_{ij}-A_{ij})\geq t\right\}$
$\leq 2\exp\left(-\frac{2t^{2}}{n}\right).$
Setting $t=\sqrt{2n\log n}$, we obtain
$\operatorname{\mathbb{P}}\{|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\geq\sqrt{2n\log n}\}\leq 2e^{-4\log n}=\frac{2}{n^{4}},$ (47)
which establishes (31).
For (32), by the mean value theorem, we have
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|=\left|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})\right|\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j})\right|,$
for some $c_{j}\in[d_{j}(\mathbf{R})-|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|,d_{j}(\mathbf{R})+|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|]$.
Using the bound from (47), it follows that
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|\leq\sqrt{2n\log n}\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|,$
for some $c_{j,1}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n},d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
$\blacksquare$
Lemma 6.
Under Assumptions (1)-(2),
$\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\leq C\sqrt{n\log n},$ (48)
with probability at least $1-\frac{14}{n^{2}}$, where $C$ is a fixed positive constant independent of $n$.
Proof of Lemma 6:
Since both $\mathbf{D}_{\mathbf{A}}$ and $\mathbf{D}_{\mathbf{R}}$ are diagonal matrices, we first derive an entrywise bound. For any $i=1,\ldots,n$, we have

### v3

@@ 4.2 Proof of the Second Part of Theorem 1
(B.26) $|\lambda_{2}(\mathbf{T})-\lambda_{2}(\mathbf{T}^{*})|\leq\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert.$
Therefore, it suffices to bound the spectral norm, $\|\cdot\|$, on the right-hand side.
Furthermore, under Assumptions 1 and 2 we have
$0<\liminf_{n}\frac{\min_{i}d_{i}(\mathbf{R})}{\max_{i}d_{i}(\mathbf{R})}\leq\limsup_{n}\frac{\max_{i}d_{i}(\mathbf{R})}{\min_{i}d_{i}(\mathbf{R})}<\infty,$
to ensure that $d_{j}(\mathbf{R})/d_{i}(\mathbf{R})$ is bounded by positive constants $K_{1},K_{2}$ for any $i,j$. Combining with the properties of $\phi$ (as utilized in Lemma B.4, Appendix 4.1), it follows that the ratio $\phi\left(\alpha,d_{j}(\mathbf{R})\right)/\phi\left(\alpha,d_{i}(\mathbf{R})\right)$ is bounded by positive constants $C_{\phi,\min}$ and $C_{\phi,\max}$ (which depend on $\alpha,K_{1},K_{2}$, but not on $n$ or $d_{i}(\mathbf{R})$ itself). Thus, for each $i$:
$\left(\mathbf{D}_{\mathbf{R}}\right)_{ii}=\sum_{j}R_{ij}\frac{\phi\left(\alpha,d_{j}(\mathbf{R})\right)}{\phi\left(\alpha,d_{i}(\mathbf{R})\right)}\geq\left(\sum_{j}R_{ij}\right)C_{\phi,\min}=d_{i}(\mathbf{R})\cdot C_{\phi,\min}.$
Therefore, we have, for some positive constant $C_{\phi,\min}$ that is independent of $n$,
$\min_{i}\left(\mathbf{D}_{\mathbf{R}}\right)_{ii}\geq C_{\phi,\min}\cdot\min_{i}d_{i}(\mathbf{R}).$
Then, note that the spectral norm in (B.26) can be decomposed as:
(B.27) $\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert$
@@
(B.28) $\|\mathbf{A}\|\leq\|\mathbf{A}-\mathbf{R}\|+\|\mathbf{R}\|\,.$
@@
(B.29) $\|\mathbf{R}\|\leq\|\mathbf{R}\|_{F}=\sqrt{\sum_{i,j}R^{2}_{ij}}\leq Cn\max_{i,j}p_{ij}\leq C\min_{j}d_{j}(\mathbf{R})\,$,
@@
Applying Proposition B.1 along with Assumption 1, we obtain
(B.30) $\|\mathbf{A}\|\leq C\min_{j}d_{j}(\mathbf{R})\,$,
with probability at least $1-{2}/{n^{2}}$. Similarly, by Proposition B.2,
(B.31) $\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\right\rVert\leq C\left(\min_{j}d_{j}(\mathbf{R})\right)^{-1/2}\,$,
with probability at least $1-{14}/{n^{2}}$, for sufficiently large $n$.
@@
(B.32) $\left\{\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\right\}\bigcap\left\{\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\right\},$
which holds with probability at least $1-{16}/{n^{2}}$, we derive
(B.33) $\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert\leq C\frac{\sqrt{n\log n}}{\min_{i}d_{i}(\mathbf{R})}=C\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,.$
@@
Proof of Remark 1
The proof is similar to Proposition 1 and Theorem 1. The main difference lies in controlling $|d_{i}^{\alpha_{i}}-d_{i}^{\alpha_{i}}(\mathbf{R})|$ for varying $\alpha_{i}$. Since $\max_{i,j}|\alpha_{i}-\alpha_{j}|\leq c$, the values of $\alpha_{i}$ concentrate around $\bar{\alpha}=\sum_{i=1}^{n}\alpha_{i}/n$, implying that $|d_{i}^{\alpha_{i}}-d_{i}^{\alpha_{i}}(\mathbf{R})|$ is of the same order as $|d_{i}^{\bar{\alpha}}-d_{i}^{\bar{\alpha}}(\mathbf{R})|$. $\Box$
Appendix C Appendix C. Proofs of Results in Section 4
1 Wisdom: Related Proofs
@@
Given any two-type model and the corresponding expected network $\mathbf{R}$, let $d_{\theta\theta^{\prime}}$ be the expected number of type-$\theta^{\prime}$ agents linked to each type-$\theta$ agent, $\theta,\theta^{\prime}\in\{h,l\}$.
Recall that there are $n_{\theta}$ type-$\theta$ agents. By Equation (5) in Proposition 1, a type-$\theta$ agent’s influence is
$s_{\theta}=\frac{S_{\theta}}{n_{h}S_{h}+n_{l}S_{l}},$
@@
$S_{h}=d_{hh}d_{h}^{\alpha}d_{h}^{\alpha}+d_{hl}d_{h}^{\alpha}d_{l}^{\alpha},\quad\text{and}\quad S_{l}=d_{ll}d_{l}^{\alpha}d_{l}^{\alpha}+d_{lh}d_{h}^{\alpha}d_{l}^{\alpha}.$
It is easy to see that
$\frac{S_{h}}{S_{l}}=\frac{d_{hh}d_{h}^{\alpha}d_{h}^{\alpha}+d_{hl}d_{h}^{\alpha}d_{l}^{\alpha}}{d_{ll}d_{l}^{\alpha}d_{l}^{\alpha}+d_{lh}d_{h}^{\alpha}d_{l}^{\alpha}}=\frac{d_{hh}(d_{h}/d_{l})^{\alpha}+d_{hl}}{d_{ll}(d_{h}/d_{l})^{-\alpha}+d_{lh}}.$
Since $d_{h}>d_{l}$ (by definition) and $d_{hh},d_{hl},d_{lh},d_{ll}>0$ are constant in $\alpha$, we have ${S_{h}}/{S_{l}}$ increases in $\alpha$, diverges to $\infty$ as $\alpha\rightarrow+\infty$, and converges to 0 as $\alpha\rightarrow$-$\infty$.
Thus,
$s_{h}=\frac{S_{h}}{n_{h}S_{h}+n_{l}S_{l}}=\frac{1}{n_{h}+n_{l}(S_{h}/S_{l})^{-1}}$
increases in $\alpha$, and
$s_{l}=\frac{S_{l}}{n_{h}S_{h}+n_{l}S_{l}}=\frac{1}{n_{h}(S_{h}/S_{l})+n_{l}}$
decreases in $\alpha$. Finally, $s_{h}(-\infty)=0$, $s_{l}(-\infty)=1/n_{l}$, $s_{h}(+\infty)=1/n_{h}$, and $s_{l}(+\infty)=0$. $\Box$
@@
The result follows immediately from Corollary 1 and Proposition 2. $\Box$
2 Convergence Speed: Explicit Characterization and Proofs
Recall that in Section 4.2 we assume that there are $m$ groups such that $n_{1}\neq n_{2}=\ldots=n_{m}$, with linking probabilities $p_{kk}=p$ and $p_{kl}=q$, where $p\neq q$. The expected degrees are then given by $d_{1}^{*}=n_{1}p+(m-1)n_{2}q$ and $d_{2}^{*}=n_{1}q+n_{2}p+(m-2)n_{2}q$. For any weighting function $\phi$ that satisfies Properties 13, let $g(\alpha)$ denote the relative weight, as follows:
$g(\alpha)=\frac{\phi(\alpha,d_{2}^{*})}{\phi(\alpha,d_{1}^{*})}.$
We obtain the following explicit characterization of convergence speed, which immediately implies Propositions 4 and 5 in the main text (Section 4.2).
Lemma C.1
The speed of convergence, negatively related to $|\lambda_{2}|$, is as follows.
  • When $m=2$ ($m_{h}=m_{l}=1$),
    (C.1) $|\lambda_{2}(\mathbf{T}^{*})|=\left|\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\phi(\alpha,d_{2}^{*})}\right|.$
  • When $m\geq 3$ and $d_{1}^{*}<d_{2}^{*}$ ($m_{h}>1,m_{l}=1$),
@@ 2 Convergence Speed: Explicit Characterization and Proofs
    (C.2) $|\lambda_{2}(\mathbf{T}^{*})|=\begin{cases}\left|\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha\leq g^{-1}\!\left(\tfrac{n_{1}}{n_{2}}\right),\\[4.30554pt] \left|\frac{n_{2}(p-q)\,\phi(\alpha,d_{2}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha>g^{-1}\!\left(\tfrac{n_{1}}{n_{2}}\right),\end{cases}$
  • When $m\geq 3$ and $d_{1}^{*}>d_{2}^{*}$ ($m_{h}=1,m_{l}>1$),